OpenStudy (help!!!!):

2. If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together? A. 2 hours and 24 minutes B. 3 hours and 12 minutes C. 3 hours and 44 minutes D. 4 hours and 10 minutes E. 4 hours and 33 minutes

OpenStudy (anonymous):

1/t = 1/4 + 1/6 Can you continue?

OpenStudy (help!!!!):

can you show me how you got your answer

OpenStudy (help!!!!):

can you show me how you got your answer

OpenStudy (help!!!!):

that is where i am stuck on

OpenStudy (anonymous):

The common denom is 12t

OpenStudy (help!!!!):

3/12+2/12=5/12

OpenStudy (help!!!!):

then what do you do?

OpenStudy (anonymous):

The common denom is 12t, you missed t !

OpenStudy (help!!!!):

?

OpenStudy (anonymous):

How can you solve for t, without t in your equation???

OpenStudy (help!!!!):

5/12

OpenStudy (help!!!!):

then whatis next? plz?

OpenStudy (anonymous):

1/t = 1/4 + 1/6 12 = 3t + 2t -> t = 12/5 hr = 2.4 hrs = 2hr 24'

OpenStudy (help!!!!):

wait how did you get 12 = 3t + 2t or step 2?

OpenStudy (anonymous):

1/t = 1/4 + 1/6 12* 1/12 = 3t* 1/12 + 2t * 1 / 12

OpenStudy (help!!!!):

can you show me the steps